The volume of a container on board a C-5 is 405 cubic feet. The height of the container is 9 feet, and its length is 4 feet longer than the width. Starting with the equation Volume = Length x Width x Height, use the given information to set up and solve a quadratic equation to determine the length and width of the container.
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = x + 4
Height = 9
Volume = Length x Width x Height
405 = (x + 4) * (x) * 9 (* = times)
405 = 9x^2 + 36x
Put in ax^2 + bx + c form:
9x^2 + 36x - 405 = 0
Divide throughout by 9
9(x^2 + 4x - 45) = 0
Factorize
9(x - 5)(x + 9) = 0
x + 9 = 0
x = -9 (Disregard as -ve)
x - 5 = 0
x = 5
.......
Height = 9 feet
Width = 5 feet
Length = 9 feet
Hope this helps :-)